​Order: Infuse 1 mL of normal saline IV for every 2 ml of urine in excess of 50 mL. The​ client's urine output is 80 mL. How much normal saline will you​ administer?
Answer #1

For every 2 ml in excess of 50 ml,1 ml saline is administered .

80 ml= 80-50 = 30 ml excess

So amount of normal saline to be administered= 30/2 =15 ml

So 15 ml of normal saline has to be administered.